OpenStudy (anonymous):

An escalator that has 30 steps moves upward at the rate of one step per second. Franklin gets on and walks up at his own speed. At the top, he realizes that he dropped a penny on the lower floor and believing that "a penny saved is a penny earned" decides to walk back down and retrieve it. He uses the same escalator so that he can keep the penny in view. Franklin walks down at the same speed that he walked up. It takes him twice as long to get down as it took him to go up. How fast is his walking speed?

OpenStudy (experimentx):

what do you mean?? first he takes escalator and next time he walks down??

OpenStudy (anonymous):

no he takes the escalator both times...something called relative velocity has to be used i think

OpenStudy (experimentx):

speed and velocity are two different things, but i don't see any real question here. at first you see, that Franklin has velocity of 1 step per second while coming up. and while going down he has velocity of -0.5 step per second. I don't see any problem here.

OpenStudy (anonymous):

Let W be the speed he walks at (ignoring the effects of the escalator), E be the speed of the escalator, U be the total speed he moves upward at, and D be the total speed he moves downward at. Your equations are U=W+E D=W-E D=1/2U Solve for W